# Orbital Altermagnetism

### v1

@@ D.1 Definition and symmetry constraints of 2D orbital altermagnetism
We define a 2D orbital altermagnet as a system whose real space can be partitioned into two sublattices carrying nonzero but opposite orbital magnetic moments. The sublattices may consist either of atomic sites or of loop currents on bonds. Nonmagnetic systems or systems belonging to type-II magnetic space groups are excluded from this definition, since in such cases the orbital magnetization $\bm{M}({\bm{r}})$ is symmetry-forced to vanish.
For orbital altermagnetism to occur, the two sublattices must be related by at least one symmetry operation, and $L_{z}({\bm{k}})$ must not be symmetry-forced to vanish. In two dimensions, $L_{z}({\bm{k}})=0$ is enforced when either $\mathcal{PT}$ or $C_{2z}\mathcal{T}$ symmetry is present. By contrast, symmetries such as $\tau\mathcal{T}$ and $m_{z}\mathcal{T}$ are compatible with orbital altermagnetism, as they only require $L_{z}$ to be odd in momentum.
This situation differs from conventional 2D spin altermagnets without spin-orbit coupling (SOC) or complex hoppings, where these symmetries are absent. Because the spin point group element $S_{2}=\{\mathcal{T}U_{n}(\pi)|1|0\}$ (with $n$ denoting an in-plane axis perpendicular to the spin) enforces $\bm{S}({\bm{k}})=\bm{S}(-{\bm{k}})$, whereas $\tau\mathcal{T}$ and $m_{z}\mathcal{T}$ require $\bm{S}({\bm{k}})=-\bm{S}(-{\bm{k}})$. Therefore, they effectively eliminate spin splitting. When SOC is introduced, the symmetry element $S_{2}$ is no longer a symmetry of the system, removing the constraint and enabling the formation of a $p$-wave altermagnetism. However, in 2D systems with $\mathcal{P}$ or $C_{2z}$ symmetry, it remains necessary to break $\tau\mathcal{T}$ and $m_{z}\mathcal{T}$ to stabilize orbital altermagnetism.
D.2 Magnetic point groups allowing the 2D orbital-altermagnetic ferromagnet
@@
To identify the magnetic point groups that permit 2D orbital altermagnetic ferromagnets, we impose the following criteria: (i) the group must be compatible with ferromagnetism; (ii) it must contain at least one symmetry operation that reverses the $z$-component of magnetization (see Fig. 4); (iii) both $\mathcal{PT}$ and $C_{2z}\mathcal{T}$ symmetries must be excluded. Applying these screening criteria yields only six compatible magnetic point groups: $2$, $m$, $m_{z}^{\prime}$, $2/m$, $m_{z}^{\prime}m2^{\prime}$, and $m^{\prime}m_{z}^{\prime}2$. Among these, only the $2/m$ group exhibits even-parity altermagnetism, which is invariant under ${\bm{k}}\to-{\bm{k}}$.
@@
Table 1: Magnetic point groups compatible with 2D orbital altermagnetism. The criteria are: (i) compatibility with ferromagnetism, (ii) presence of at least one operation reversing the $z$-component of magnetization, and (iii) exclusion of both $\mathcal{PT}$ and $C_{2z}\mathcal{T}$ symmetries. The parity column indicates the parity under inversion ${\bm{k}}\to-{\bm{k}}$.
@@ 22 $C_{2z}$ Odd
D.3 Importance of $S_{2}$ breaking for orbital altermagnetism in collinear structures
In the absence of SOC or complex hoppings, the appropriate framework to describe the symmetry of magnetic systems is the spin group, rather than the magnetic point group. For a collinear magnetic structure with the spin aligned along the $z$-axis, two types of symmetry operations leave the spin vector $\bm{S}$ invariant [117]:
$S_{1}=\{U_{z}(\theta)|1|\bm{0}\},\,\theta\in[0,2\pi);\quad S_{2}=\{\mathcal{T}U_{xy}(\pi)|1|\bm{0}\}.$ (15)
Here, $S_{1}$ represents a rotation by angle $\theta$ about the $z$-axis, while $S_{2}$ represents a rotation $\pi$ about an in-plane axis accompanied by time reversal $\mathcal{T}$.
The orbital magnetic moment satisfies the following equation [118]:
$M_{\mathrm{orb}}({\bm{r}})=\theta\det(R)R\cdot M_{\mathrm{orb}}(\{R|t\}^{-1}{\bm{r}}),$ (16)
where $\theta=-1$ if time-reversal $\mathcal{T}$ is included in the operation, and $\theta=+1$ otherwise. $R$ represents a proper or improper rotation of the system, $t$ represents a space translation, and $\det(R)$ is the determinant of the matrix $R$. Since the real-space parts of $S_{1}$ and $S_{2}$ are identical but differ in the spin-space operations ($S_{2}$ includes time reversal $\mathcal{T}$ while $S_{1}$ does not), one finds that $M_{\mathrm{orb}}({\bm{r}})=-M_{\mathrm{orb}}({\bm{r}})=0$. Therefore, in a collinear system without SOC or complex hoppings, the orbital magnetic moment is symmetry-forbidden. In this sense, the operation $S_{2}$ acts as an effective time-reversal symmetry for the orbital sector.
In momentum space, $S_{2}$ enforces that the spin is even in momentum $\bm{S}(-{\bm{k}})=\bm{S}({\bm{k}})$, while the orbital angular momentum is odd $L_{z}(-{\bm{k}})=-L_{z}({\bm{k}})$. If the systems additionally respects either $C_{2z}$ or inversion $\mathcal{P}$ symmetry, which require $L_{z}({\bm{k}})=L_{z}(-{\bm{k}})$, then the orbital angular momentum must vanish identically: $L_{z}({\bm{k}})=0$.
By contrast, the inclusion of SOC or complex hopping explicitly breaks the $S_{2}$ symmetry. This breaking removes the effective orbital time-reversal constraint, thereby allowing finite orbital magnetic moments $M_{\mathrm{orb}}({\bm{r}})$ and nonvanishing orbital angular momentum $L_{z}({\bm{k}})$. This highlights that SOC or complex hoppings are essential for stabilizing orbital altermagnetism in collinear 2D systems.
@@ Appendix E First-Principles Calculation Method
First-principles calculations are performed within the framework of density-functional theory (DFT) using the projector augmented-wave method and the Perdew-Burke-Ernzerhof-type exchange-correlation functional [112], as implemented in the Vienna ab initio simulation package (VASP) [113]. We set the cutoff energy of the plane-wave basis to be 520 eV and adapted a $\times15$$\times$1 $\Gamma$-center k-mesh for the BZ sampling. Spin-orbit coupling (SOC) is included self-consistently in the calculations. To eliminate the spurious interaction between the 2D material and its images, a vacuum region of more than 12 Å along the $z$ direction is set in the calculation. All crystal structures are fully optimized until the forces are less than $10^{-3}$ eV/Å and the change of the total energy is less than 0.01 meV. The correlation effect for the Cr 3dd electrons and Mo 4dd electrons is treated by the DFT+$U$ method with $U=3.55$ and $2.4$ eV [114, 115], respectively. To study the band geometric quantities such as the Berry curvature and orbital magnetization, we construct maximally localized Wannier functions (MLWFs) and derive a Wannier-based ab initio tight-binding model for each structure using the Wannier90 code [116]. Based on the Wannier representation, we calculate the Berry curvature $\Omega({\bm{k}})$ and orbital magnetization of all states.
@@ F.1 Current-induced orbital magnetization in MoO monolayer
Spin–orbit coupling (SOC) opens a gap at the Weyl points $W_{1}$ and $W_{2}$ [92]. However, in the vicinity of these nearly degenerate points, the orbital angular momentum $L_{z}$ remains strongly enhanced. Under an external field, these points contribute much more significantly to orbital magnetization than spin polarization. We calculated the nonlinear current-induced spin polarization (CISP) and current-induced orbital polarization (CIOP) in the MoO monolayer, as illustrated in Fig. 5(a). The results reveal that the orbital polarization is substantially stronger than the spin polarization. As shown in Fig. 5(b), $W_{1}$ and $W_{2}$ points represent the hotspot regions in the $\bm{k}$}-space distribution of $\chi_{zyy}({\bm{k}})$. While the overall distribution follows the symmetry of the system, the local behaviors around $W_{1}$ and $W_{2}$ are quite different, as shown in Fig. 5. In particular, although the $L_{z}$ distribution is linked by $C_{4z}\mathcal{T}$, the velocity matrix elements $v_{y}$ at $W_{1}$ and $W_{2}$ lack such a symmetry connection, preventing cancellation between their contributions.

### v2

@@ D.1 Definition and symmetry constraints of 2D orbital altermagnetism
We define a 2D orbital altermagnet as a system whose real space can be partitioned into two sublattices carrying nonzero but opposite orbital magnetic moments. The sublattices may consist either of atomic sites or of loop currents on bonds. In both cases, we focus on the out-of-plane component $M_{z}({\bm{r}})$, which is the natural component in a two-dimensional geometry.
A key feature of orbital altermagnetism is that the net out-of-plane orbital magnetization vanishes, which the local distribution is not identically zero. This is the orbital analogue of a compensated antiferromagnetic-like or altermagnetic-like pattern, with the magnetic degree of freedom originating from orbital motion rather than from spin.
In momentum space, for an orbital altermagnet, the momentum-resolved orbital magnetic moment or the orbital angular momentum $L_{z}({\bm{k}})$ must not be forced to vanish by symmetry across the Brillouin zone. In particular, any antiunitary symmetry that maps ${\bm{k}}\to{\bm{k}}$ while flipping the sign of the axial out-of-plane moment would enforce $L_{z}({\bm{k}})=0$, and is therefore incompatible with a finite orbital moment texture. This includes, for example, $\mathcal{PT}$ or $C_{2z}\mathcal{T}$.
D.2 Screening magnetic point groups compatible with 2D orbital altermagnetism
We performed an exhaustive symmetry screening over all magnetic point groups to identify those that can host 2D orbital altermagnetism. The screening procedure follows directly from the definition and the symmetry constraints discussed above.
(i) Remark on gray magnetic point groups and the underlying MSG type.
A gray (type-II) magnetic point group contains the pure time-reversal operation $\mathcal{T}$ at the point-group level. Importantly, the same gray point group can originate from two physically different situations at the magnetic space-group (MSG) level: a type-II (gray) MSG, where $\mathcal{T}$ appears as $\{1|\mathbf{0}\}\mathcal{T}$ and the crystal is nonmagnetic; or a type-IV MSG, where the relevant antiunitary symmetry is a generalized time reversal of the form
$\tilde{\mathcal{T}}=\{1|\boldsymbol{\tau}\}\mathcal{T},$ (17)
with a nontrivial translation $\boldsymbol{\tau}$, which projects to $\mathcal{T}$ in the point-group description.
Since the definition of 2D orbital altermagnetism concerns magnetic states with a nontrivial local orbital-moment pattern, we do not consider the nonmagnetic realization (type-II MSG) of gray point groups. Therefore, throughout the screening below, whenever a gray (type-II) magnetic point group is encountered, it is implicitly understood as the realization compatible with a type-IV MSG (i.e., with $\tilde{\mathcal{T}}=\{1|\boldsymbol{\tau}\}\mathcal{T}$ rather than $\{1|\mathbf{0}\}\mathcal{T}$).
(ii) Only operations preserving the $k_{z}=0$ plane.
Since we target a two-dimensional orbital texture characterized on the $k_{z}=0$ plane, we only keep symmetry operations that leave this plane invariant. Concretely, we consider the subgroup of operations $g$ such that
$g:\ (k_{x},k_{y},0)\mapsto(k_{x}^{\prime},k_{y}^{\prime},0).$ (18)
Operations that mix $k_{z}$ with in-plane components are discarded for the purpose of the 2D classification.
(iii) Existence of a symmetry that flips the out-of-plane orbital moment.
To realize the defining compensated pattern (zero net out-of-plane moment but nonzero local $M_{z}$), the symmetry group must contain at least one operation that maps the two sublattices onto each other while reversing the sign of the out-of-plane axial moment. In other words, we require the presence of a symmetry element $g$ such that
$M_{z}({\bm{r}})\xrightarrow{g}-\,M_{z}(g{\bm{r}}),$ (19)
which guarantees cancellation of the net $z$-component upon summation over the symmetry-related sublattices (or loop-current substructures), while still allowing $M_{z}({\bm{r}})\not\equiv 0$.
(iv) Exclusion of symmetries that force $L_{z}({\bm{k}})$ to vanish.
Finally, we exclude magnetic point groups containing any (antiunitary) operation that enforces the constraint
$L_{z}({\bm{k}})=-L_{z}({\bm{k}}),$ (20)
which would symmetry-force $L_{z}({\bm{k}})\equiv 0$ on the $k_{z}=0$ plane and thus forbid a finite orbital texture. In practice, this removes cases where an antiunitary symmetry maps ${\bm{k}}\to{\bm{k}}$ while flipping the sign of the axial out-of-plane moment (for example, $\mathcal{P}\mathcal{T}$ or $C_{2z}\mathcal{T}$, as discussed above).
(v) Conventions and coordinate choices.
Throughout this work, symmetry operations are taken in their standard settings with conventional axis orientations. For a given material, the crystallographic axes may be rotated relative to these standard conventions, and a corresponding change of coordinates may be required to apply the screening rules to a specific structure. Such material-dependent axis redefinitions do not change the symmetry logic, but they can affect which component is identified as the out-of-plane direction in the chosen coordinate system. A case-by-case coordinate alignment is therefore left for future, system-specific analyses and is not pursued in detail here. For example, consider the magnetic point group $m^{\prime}m2^{\prime}$. In the standard convention one may represent its generators as $m_{x}\mathcal{T}$, $m_{y}$, and $C_{2z}\mathcal{T}$; in this orientation $C_{2z}\mathcal{T}$ can map ${\bm{k}}\!\to\!{\bm{k}}$ while reversing the axial out-of-plane moment, thereby symmetry-forcing $L_{z}({\bm{k}})$ to vanish and excluding 2D orbital altermagnetism. However, under a different (rotated) axis assignment for the same abstract group, one may instead take $m_{z}\mathcal{T}$, $m_{x}$, and $C_{2y}$ as generators; in that coordinate choice, the corresponding constraints no longer forbid a finite $L_{z}({\bm{k}})$ texture, and 2D orbital altermagnetism becomes symmetry-allowed.
Applying the above criteria yields a restricted set of magnetic point groups compatible with two-dimensional orbital altermagnetism, which we summarize in Table 1 and discuss below.
Table 1: Magnetic point groups (MPGs) that allow 2D orbital altermagnetism, together with the symmetry-enforced leading splitting wave of the orbital texture. The classification is given in the standard setting with conventional axis orientations; for a specific material, an appropriate rotation of the coordinate axes may be required.
Leading wave Allowed magnetic point groups (MPGs)
p-wave $\mathrm{1.1^{\prime},\ 2.1,\ 2.1^{\prime},\ m.1,\ m.1^{\prime}}$
d-wave $\mathrm{2/m.1,\ 222.1,\ mm2.1,}$
$\mathrm{mmm.1,\ 4^{\prime},\ -4^{\prime},\ 4^{\prime}/m,\ 4^{\prime}22^{\prime},\ 4^{\prime}m^{\prime}m,\ -4^{\prime}2^{\prime}m,\ -4^{\prime}2m^{\prime},\ 4^{\prime}/mm^{\prime}m}$
g-wave $\mathrm{422.1,\ 4mm.1,\ -42m.1,\ 4/mmm,}$
f-wave $\mathrm{3.1^{\prime},\ 32.1,\ 32.1^{\prime},\ 3m.1,\ 3m.1^{\prime},\ -6.1^{\prime},\ -6^{\prime},\ -6m2.1,\ -6m2.1^{\prime},\ -6^{\prime}m^{\prime}2,\ -6^{\prime}m2^{\prime}}$
i-wave $\mathrm{-3m.1,\ 622.1,\ 6mm.1,\ 6/mmm.1,}$
D.3 Magnetic point groups allowing the 2D orbital-altermagnetic ferromagnet
@@
To identify the magnetic point groups that permit 2D orbital altermagnetic ferromagnets, we impose the following criteria: (i) the group must be compatible with ferromagnetism; (ii) it must contain at least one symmetry operation that reverses the $z$-component of magnetization (see Fig. 4); (iii) both $\mathcal{PT}$ and $C_{2z}\mathcal{T}$ symmetries must be excluded. Applying these screening criteria yields only six compatible magnetic point groups: $2$, $m$, $m_{z}^{\prime}$, $2/m$, $m_{z}^{\prime}m2^{\prime}$, and $m^{\prime}m_{z}^{\prime}2$. Among these, only the $2/m$ group exhibits even-parity altermagnetism, which is invariant under ${\bm{k}}\to-{\bm{k}}$ (see Table 2).
@@
Table 2: Magnetic point groups allowing 2D orbital altermagneic ferromagnets. The criteria are: (i) compatibility with ferromagnetism, (ii) presence of at least one operation reversing the $z$-component of magnetization, and (iii) exclusion of both $\mathcal{PT}$ and $C_{2z}\mathcal{T}$ symmetries. The parity column indicates the parity under inversion ${\bm{k}}\to-{\bm{k}}$.
@@ 22 $C_{2y}$ Odd
D.4 Role of spin-orbit coupling in 2D orbital altermagnetism
We emphasize that spin–orbit coupling (SOC) is central for the two-dimensional setting considered in this work. In the absence of SOC, spin and orbital sectors are decoupled. As a result, symmetry operations acting in spin space influence the orbital magnetic moment only through whether the associated symmetry is unitary or antiunitary when projected onto the orbital sector. In particular, for collinear or coplanar magnetic configurations, one can define an effective orbital time-reversal-like antiunitary symmetry, which enforces the local orbital moment in real space to vanish identically. Therefore, such SOC-free collinear/coplanar states do not satisfy our definition of orbital altermagnetism, which requires $M_{\rm orb}({\bm{r}})\not\equiv 0$.
More specifically, the real-space orbital magnetic moment $M_{\rm orb}({\bm{r}})$ (an axial vector) transforms under a general spin space-group operation $\{W|R|\mathbf{t}\}$ as
$M_{\rm orb}({\bm{r}})=\det(W)\,\det(R)\,R\cdot M_{\rm orb}\!\left(\{R|\mathbf{t}\}^{-1}{\bm{r}}\right),$ (21)
where $\det(W)=-1$ can be viewed as an effective orbital time-reversal symmetry, $R$ is a proper or improper rotation, and $\mathbf{t}$ is a translation. Equation (21) implies that if the symmetry group contains an antiunitary operation of the form $\{W|1|\mathbf{0}\}$ with $\det(W)=-1$ that leaves $\bm{r}$} invariant, then
$M_{\rm orb}({\bm{r}})=-\,M_{\rm orb}({\bm{r}})=0,$ (22)
so the local orbital moment is symmetry-forced to vanish everywhere. Such operations can be viewed as an “orbital time-reversal” symmetry in the SOC-free limit  [128].
For collinear magnetic configurations with spins aligned along the $z$ axis, the effective orbital time-reversal symmetry can be written as $\{\mathcal{T}U_{xy}(\pi)|1|\mathbf{0}\}$, consisting of time reversal followed by a $\pi$ spin rotation about an arbitrary in-plane axis. For coplanar structures with spins lying in the $xy$ plane, the corresponding effective orbital time-reversal symmetry is $\{\mathcal{T}U_{z}(\pi)|1|\mathbf{0}\}$, which transforms a spin moment $(S_{x},S_{y},S_{z})$ to $(S_{x},S_{y},-S_{z})$. Consequently, without SOC, orbital altermagnetism is only possible for noncoplanar magnetic structures, where no such effective orbital time-reversal symmetry exists.
Finally, we note an additional simplification in the SOC-free limit. Because spin-space operations $W$ affect the orbital magnetic moment only through whether $\det(W)=+1$ or 1-1, their action on $M_{\rm orb}$ reduces to a binary classification rather than depending on the detailed spin rotation itself. As a consequence, any nontrivial spin point group can be associated with a corresponding magnetic point group that captures the induced constraints on the orbital sector (see Table 7 in  [129]). Therefore, when SOC is neglected, one may determine whether an orbital altermagnetism is symmetry-allowed by identifying the magnetic point group corresponding to the noncoplanar configuration’s nontrivial spin point group and then applying our magnetic-point-group screening results.
@@ Appendix E First-Principles Calculation Method
First-principles calculations are performed within the framework of density-functional theory (DFT) using the projector augmented-wave method and the Perdew-Burke-Ernzerhof-type exchange-correlation functional [118], as implemented in the Vienna ab initio simulation package (VASP) [119]. We set the cutoff energy of the plane-wave basis to be 520 eV and adapted a $\times15$$\times$1 $\Gamma$-center k-mesh for the BZ sampling. SOC is included self-consistently in the calculations. To eliminate the spurious interaction between the 2D material and its images, a vacuum region of more than 12 Å along the $z$ direction is set in the calculation. All crystal structures are fully optimized until the forces are less than $10^{-3}$ eV/Å and the change of the total energy is less than 0.01 meV. The correlation effect for the Cr 3dd electrons and Mo 4dd electrons is treated by the DFT+$U$ method with $U=3.55$ and $2.4$ eV [120, 121], respectively. To study the in-plane ferromagnetic configuration, we further performed constrained-moment calculations in VASP using the tags I_CONSTRAINED_M and M_CONSTR to align the spin moments along a chosen in-plane direction. To study the band geometric quantities such as the Berry curvature and orbital magnetization, we construct maximally localized Wannier functions (MLWFs) and derive a Wannier-based ab initio tight-binding model for each structure using the Wannier90 code [122]. Based on the Wannier representation, we calculate the Berry curvature $\Omega({\bm{k}})$ and orbital magnetization of all states. All input structures, calculation parameters, and Wannier90 projectors used in this work are openly available in the Zenodo repository at 10.5281/zenodo.17506196.
@@ Appendix F Numerical results of other candidate materials for orbital altermagnetism
Figure 5: Momentum-space distributions of quantities related to the orbital angular moment for the highest valence band ($n=v$) and the lowest conduction band ($m=c$) of the MoO monolayer. (a) Distribution of the total $L_{z,v}$ in the Brillouin zone. (b) Distribution of the interband-resolved contribution $L_{z,v}^{(c)}$. (c) Distribution of $\textrm{Im}[v_{x}^{vc},v_{y}^{cv}]$, showing the nodal lines separating regions of opposite sign. (d) Distribution of the Berry curvature $\Omega_{z,v}$ for comparison.
@@ F.1 Current-induced orbital magnetization in MoO monolayer
We first present the $\bm{{\bm{k}}}$-space distribution of quantities related to the orbital angular moment. Figure 5(b) shows the $L_{z}$ of the highest valence band $v$ mainly originates from the lowest conduction band $c$. Here we define:
$\begin{split}&M_{n{\bm{k}}}^{(m)}=-\frac{e}{2\hbar c}\,\mathrm{Im}\frac{\langle u_{n{\bm{k}}}|\hat{\mathbf{v}}|u_{m{\bm{k}}}\rangle\times\langle u_{m{\bm{k}}}|\hat{\mathbf{v}}|u_{n{\bm{k}}}\rangle}{\varepsilon_{n{\bm{k}}}-\varepsilon_{m{\bm{k}}}}\qquad(m\neq n),\\ &L_{z,n}^{(m)}({\bm{k}})=-g_{L}\mu_{B}[M_{n{\bm{k}}}^{(m)}]_{z}.\end{split}$ (23)
To investigate the origin of the frequent sign changes of $L_{z}$ in k-space, we focus on the numerator in the expression Im[$v_{x}^{vc}$$v_{x}^{vc}$]=Im($v_{x}^{vc}$$v_{y}^{cv}$-$v_{y}^{vc}$$v_{x}^{cv}$), and plot its $\bm{k}$}-space distribution for the $c$ and $v$ bands, as shown in Fig. 5(c). In momentum space, Im[$v_{x}^{vc}$$v_{x}^{vc}$] typically changes sign across the lines or regions where it vanishes. These zero-value contours divide the Brillouin zone into several sectors, and crossing such a line naturally leads to a reversal of the sign of the orbital magnetic moment. Similarly, the numerator of the Berry curvature has a comparable form. Therefore, we also plotted the Berry curvature distribution in $\bm{k}$}-space, which shows a similar sign-reversal behavior.
Spin–orbit coupling opens a gap at the Weyl points $W_{1}$ and $W_{2}$ [95]. However, in the vicinity of these nearly degenerate points, the orbital angular momentum $L_{z}$ remains strongly enhanced. Under an external field, these points contribute much more significantly to orbital magnetization than spin polarization. We calculated the nonlinear current-induced spin polarization (CISP) and current-induced orbital polarization (CIOP) in the MoO monolayer, as illustrated in Fig. 6(a). The results reveal that the orbital polarization is substantially stronger than the spin polarization. As shown in Fig. 6(b), $W_{1}$ and $W_{2}$ points represent the hotspot regions in the $\bm{k}$}-space distribution of $\chi_{zyy}({\bm{k}})$. While the overall distribution follows the symmetry of the system, the local behaviors around $W_{1}$ and $W_{2}$ are quite different, as shown in Fig. 6. In particular, although the $L_{z}$ distribution is linked by $C_{4z}\mathcal{T}$, the velocity matrix elements $v_{y}$ at $W_{1}$ and $W_{2}$ lack such a symmetry connection, preventing cancellation between their contributions.